# The Quantum Measurement Problem: A Review of Recent Trends

### v2

@@ 5.5 Deterministic Models
Bell’s theorem shows that no local hidden variable theories can reproduce all quantum mechanical predictions [104]. Thus, the experimental violation of Bell’s inequalities seemed to be the end of any local deterministic model. However, Bell’s theorem explicitly requires the validity of statistical independence assumption, which states that the experimental settings are not influenced by any hidden factors determining the system’s evolution. This section surveys two local deterministic models—Superdeterminism and Retrocausality—built on the challenge of the statistical independence assumption.
Superdeterminism posits that hidden correlations between measurement settings and system properties predetermine quantum outcomes [123]. For instance, a recent proposal by Donadi and Hossenfelder [124] introduced a local, deterministic model for wave function collapse that explicitly violates statistical independence while reproducing standard quantum predictions. Their model suggests that all quantum states, except measurement eigenstates, are unstable under hidden-variable perturbations and deterministically collapse before reaching the detector. Unlike objective collapse models (see Section 5.3), the collapse dynamics in the Donadi-Hossenfelder model are entirely deterministic.
@@
Retrocausal Interpretations also challenge the statistical independence assumption, but, even bolder than Superdeterminism, it assumes influence from future events. Such interpretations propose that the Measurement Problem arises from assuming a one-way flow of time. Models such as the two-state vector formalism [126] describe quantum systems by an initial state evolving forward in time and a final state evolving backward from the measurement outcome. The interaction of these forward- and backward-evolving states determines the probabilities of different results. The Pechukas’ forces [127], which are still the main justification used today for momentum rescaling along nonadiabatic coupling vectors in molecular dynamics [128], were derived from this same type of forward and backward propagation to determine quantum transitions in atomic collisions.
Retrocausality is still an active research field. Adlam [129] proposed that accepting nonlocality and imposing relativistic constraints (such as the nonexistence of a preferred reference frame) leads to retrocausality. Sutherland [130] derives the Born rule within a retrocausal model. This is a significant result because it changes the status of the rule from a postulate to a natural consequence of incorporating time-symmetric boundary conditions in quantum mechanics. A notable development is the Fixed-Point Formulation proposed by Ridley and Adlam, which presents an atemporal, all-at-once framework for retrocausality and includes a full derivation of the Born rule [131, 132]. Another recent contribution is van der Pals’ retrocausal hidden-variable model [133], which attributes the emergence of definite outcomes to resonance conditions between advanced and retarded virtual photons. These photons arise from an underlying, time-symmetric periodic process occurring beneath the Heisenberg uncertainty threshold.
@@ 5.6.1 Classical-Apparatus Inducing Collapse
The core of the Copenhagen Interpretation of quantum mechanics, developed by Bohr and Heisenberg, is the assumption that any experiment in physics must be described in terms of classical physical concepts [39, Introduction]. When such an experiment is in the realm of quantum mechanics, the tension between the system’s quantum description and the apparatus’s classical description is the crucial factor causing statistical uncertainty and ultimately selecting a single output [135]. There is no claim that the classical apparatus is not composed of quantum particles. It is just that the description of the experiment is restricted to the use of classical terminology.
@@
In a series of works [138, 139], Schonfeld has explored cloud chambers, Geiger counters, and Stern-Gerlach experiments for empirical tests of quantum measurement, including possible deviations from the Born rule in detecting rare events. He challenges the conventional view of measurement as an axiomatic feature of quantum mechanics, proposing that measurement emerges phenomenologically from collective interactions in quantum detection systems. He analyzes the detailed microstructure of real measurement systems, using classical idealizations selectively where it seems mathematically or intuitively reasonable.
@@ 5.6.2 Mind-Inducing Collapse
Then, there is an epistemic proposal where the mind doesn’t objectively cause the collapse. Still, the collapse is just an illusion caused by the observer becoming aware of the measurement outcome. We find such a perspective in London and Bauer [73, p. 251] who, in 1939, attributed the essential role played by consciousness to be “the increase of knowledge, acquired by the observation.”
@@ 5.7 Measurement in Quantum Field Theory
A complementary approach is proposed by Polo-Gómez, Garay, and Martín-Martínez, who develop a measurement theory for QFT based on spatially smeared particle detector models—such as the Unruh-DeWitt detector—that interact locally and covariantly with quantum fields [149]. Their framework, applicable to general fields and spacetime geometries, avoids problems like Sorkin’s impossible measurements and allows localized interactions without invoking instantaneous collapse. While Bostelmann et al. pursue a mathematically rigorous algebraic route, Polo-Gómez et al. adopt a more operational strategy grounded in how detectors realistically interact with quantum fields.
@@ 6 The Ontological Status of the Wave Function
A different approach comes from Nomological Interpretations, which suggest that the wave function is not a physical object but a law of nature governing quantum behavior [152]. This is a common stance in interpreting Bohmian Mechanics [111, 160].
The Berry phase, which may arise during the cyclic evolution of quantum systems [161, 162], is a mathematical feature of the wave function that could be explored to assess its ontological status. Such a phase can build, for instance, when the nuclear geometry of a molecule evolves along a closed path encircling a conical intersection. (A conical intersection is a molecular geometry at which two adiabatic electronic states degenerate, forming a conical topology of the potential energy surfaces in nuclear coordinate space [163].) Unlike dynamic phases, the Berry phase is tied to the wave function’s geometric properties and may, in principle, manifest in observable interference effects. Valahu et al. [164] claimed to have observed such geometric-phase interference in a trapped-ion quantum simulator mimicking conical intersection dynamics, which would strongly support the -ontic interpretation. However, it’s unclear if such effects could be observed in actual molecules since their experiment could be considered more a controlled simulation than a molecular measurement. Moreover, Min et al. [165] and, more recently, Ibele et al. [166] have provided strong theoretical evidence that the Berry phase originates from the Born-Oppenheimer approximation and disappears in an exact electron-nuclear treatment, suggesting it may be an artifact rather than a fundamental quantum property.
@@
7 Conclusion
It is easy to get lost in the Quantum Measurement Problem. The gargantuan amount of literature, the innumerable distinct theories, the interminable debates, and the entanglement between physics, philosophy, and formal logic all lend anyone stepping into this field a feeling analogous to being trapped in a dense tropical forest. This review intends to be an updated global map, which, although it cannot point out the escape route, as it simply doesn’t exist, will at least help us stay far from dangerous cliffs.
Here, we tried to offer a general overview of possible solutions to the Measurement Problem, which remains actively discussed by physicists and philosophers, and of newly proposed theories that we consider particularly promising or disruptive. They fall into five classes of explanations: Many-Worlds Interpretations, Epistemic Interpretations, Objective Collapse Theories, Hidden-Variable Theories, and Dualistic Collapse Hypotheses.
Today, the measurement problem is much less puzzling than it once was. Thanks to the decoherence program, we understand how quantum interferences are suppressed and how the environment selects the outcome basis. Nevertheless, we still lack a fundamental explanation for why only a single outcome is observed in experiments, and we are still taking the first steps toward a well-defined measurement theory in Quantum Field Theory. Indeed, given that QFT is likely the most successful physical theory ever developed, any satisfactory solution for the Measurement Problem must also be an acceptable solution for measurement in QFT. Therefore, it is worrisome to note that researchers on the Measurement Problem don’t usually deal with QFT (and vice versa).
Most scientists who use quantum mechanics in their daily work focus on its practical applications rather than its interpretation. Indeed, it makes no difference if we interpret the result of an experiment as the quantum state collapsing into a single branch or experiencing a branch that cuts communication from all others. Such a subtle distinction belongs to defining a preferred worldview. It is even less relevant than choosing between Newtonian or Hamiltonian descriptions of classical dynamics, which have at least instrumental implications. The exception concerns objective collapse theories, which propose new physics beyond conventional quantum mechanics. In any case, it makes sense to keep challenging all interpretations and theories to prune the weak and reach complete consistency on the remaining, even if we never get a consensual single description of quantum mechanics.

### v3

@@ 5.5 Deterministic Models
Bell’s theorem shows that no local hidden variable theories can reproduce all quantum mechanical predictions [112]. Thus, the experimental violation of Bell’s inequalities seemed to be the end of any local deterministic model. However, Bell’s theorem explicitly requires the validity of the statistical independence assumption, which states that the experimental settings are not influenced by any hidden factors determining the system’s evolution. This section surveys two local deterministic models—Superdeterminism and Retrocausality—built on the challenge to the statistical independence assumption.
Superdeterminism posits that hidden correlations between measurement settings and system properties predetermine quantum outcomes [132]. ’t Hooft proposed a concrete realization of this idea in his Cellular Automaton Interpretation (CAI) [133]. In this model, the universe is fundamentally described by a discrete, classical system evolving in time, much like a computer program updating the values of a grid according to fixed rules. These underlying configurations, called ontological states, evolve deterministically and form the actual physical reality. The familiar quantum formalism emerges only as a statistical approximation of our ignorance about the precise ontological state. In this view, superpositions are not real: they are computational tools that summarize ensembles of possible classical states. ’t Hooft argues that violations of Bell inequalities do not imply nonlocality, but instead reflect our mistaken identification of such statistical templates with physical systems.
A recent proposal by Donadi and Hossenfelder [134] introduced a local, deterministic model for wave function collapse that explicitly violates statistical independence while reproducing standard quantum predictions. Their model suggests that all quantum states, except measurement eigenstates, are unstable under hidden-variable perturbations and deterministically collapse before reaching the detector. Unlike objective collapse models (see Section 5.3), the collapse dynamics in the Donadi-Hossenfelder model are entirely deterministic.
@@
Retrocausal Interpretations also challenge the statistical independence assumption, but, even bolder than Superdeterminism, it assumes influence from future events. Such interpretations propose that the Measurement Problem arises from assuming a one-way flow of time. Models such as the Two-State Vector Formalism [136] describe quantum systems by an initial state evolving forward in time and a final state evolving backward from the measurement outcome. The interaction of these forward- and backward-evolving states determines the probabilities of different results. The Pechukas’ forces [137], which are still the primary justification used today for momentum rescaling along nonadiabatic coupling vectors in molecular dynamics [138], were derived from this same type of forward and backward propagation to determine quantum transitions in atomic collisions.
Retrocausality is still an active research field. Adlam [139] proposed that accepting nonlocality and imposing relativistic constraints (such as the nonexistence of a preferred reference frame) leads to retrocausality. Sutherland [140] derived the Born rule within a retrocausal model. This is a significant result because it changes the status of the rule from a postulate to a natural consequence of incorporating time-symmetric boundary conditions in quantum mechanics. A notable development is the Fixed-Point Formulation proposed by Ridley and Adlam, which presents an atemporal, all-at-once framework for retrocausality and includes a full derivation of the Born rule [141, 142]. Another recent contribution is van der Pals’ retrocausal hidden-variable model [143], which attributes the emergence of definite outcomes to resonance conditions between advanced and retarded virtual photons. These photons arise from an underlying, time-symmetric periodic process occurring beneath the Heisenberg uncertainty threshold.
@@ 5.6.1 Classical-Apparatus Inducing Collapse
The core of the Copenhagen Interpretation of quantum mechanics, developed by Bohr and Heisenberg, is the assumption that any experiment in physics must be described in terms of classical physical concepts [43, Introduction]. When such an experiment falls within the realm of quantum mechanics, the tension between the system’s quantum description and the apparatus’s classical description is the crucial factor that causes statistical uncertainty and ultimately selects a single output [145]. There is no claim that the classical apparatus is not composed of quantum particles. It is just that the description of the experiment is restricted to the use of classical terminology.
@@
In a series of works [148, 149], Schonfeld has explored cloud chambers, Geiger counters, and Stern-Gerlach experiments for empirical tests of quantum measurement, including possible deviations from the Born rule in detecting rare events. He challenges the conventional view that measurement is an axiomatic feature of quantum mechanics, proposing that measurement emerges phenomenologically from collective interactions in quantum detection systems. He analyzed the detailed microstructure of real measurement systems, using classical idealizations selectively where it seems mathematically or intuitively reasonable.
@@ 5.6.2 Mind-Inducing Collapse
Then, there is an epistemic proposal where the mind does not objectively cause the collapse. Instead, the apparent collapse is an illusion arising when the observer becomes aware of the measurement outcome. We find such a perspective in London and Bauer [80, p. 251] who, in 1939, attributed the essential role played by consciousness to be “the increase of knowledge, acquired by the observation.”
@@ 5.7 Measurement in Quantum Field Theory
A complementary approach is proposed by Polo-Gómez, Garay, and Martín-Martínez, who develop a measurement theory for QFT based on spatially smeared particle detector models—such as the Unruh-DeWitt detector—that interact locally and covariantly with quantum fields [159]. Their framework, applicable to general fields and spacetime geometries, avoids problems such as Sorkin’s impossible measurements and allows localized interactions without invoking instantaneous collapse. While Bostelmann et al. pursue a mathematically rigorous algebraic route, Polo-Gómez et al. adopt a more operational strategy grounded in how detectors realistically interact with quantum fields.
@@ 6 The Ontological Status of the Wave Function
A different approach comes from the Nomological Interpretation, which suggests that the wave function is not a physical object but a law of nature governing quantum behavior [162]. This is a common stance in interpreting Bohmian Mechanics [119, 170].
The Berry phase, which may arise during the cyclic evolution of quantum systems [171, 172], is a mathematical feature of the wave function that can be explored to assess its ontological status. Such a phase can build, for instance, when the nuclear geometry of a molecule evolves along a closed path encircling a conical intersection. (A conical intersection is a molecular geometry at which two adiabatic electronic states degenerate, forming a conical topology of the potential energy surfaces in nuclear coordinate space [173].) Unlike dynamic phases, the Berry phase is tied to the wave function’s geometric properties and may, in principle, manifest in observable interference effects. Valahu et al. [174] claimed to have observed such geometric-phase interference in a trapped-ion quantum simulator mimicking conical intersection dynamics, which would strongly support the -ontic interpretation. However, it is unclear whether such effects could be observed in actual molecules since their experiment could be considered more of a controlled simulation than a molecular measurement. Moreover, Min et al. [175] and, more recently, Ibele et al. [176] have provided strong theoretical evidence that the Berry phase originates from the Born-Oppenheimer approximation and disappears in an exact electron-nuclear treatment, suggesting it may be an artifact rather than a fundamental quantum property.
@@
7 Through the Forest of Quantum Foundations
It is easy to get lost in the Measurement Problem. The gargantuan amount of literature, the innumerable distinct theories, the endless debates, and the entanglement between physics, philosophy, and formal logic all lend anyone stepping into this field a feeling analogous to being trapped in a dense tropical forest. This review aims to offer an updated global map, which, although it cannot pinpoint the escape route, as it simply doesn’t exist, will at least help us steer clear of dangerous cliffs.
Here, we provide a general overview of possible solutions to the Measurement Problem, which remains an actively discussed topic among physicists and philosophers, as well as of newly proposed theories that we consider particularly promising or disruptive. They fall into five classes of explanations: Many-Worlds Interpretations, Epistemic Interpretations, Objective Collapse Theories, Hidden-Variable Theories, and Dualistic Collapse Hypotheses.
Today, the Measurement Problem is much less puzzling than it once was. Thanks to the decoherence program, we understand how quantum interferences are suppressed and how the environment selects the outcome basis. Nevertheless, we still lack a fundamental explanation for why only a single outcome is observed in experiments, and we are still taking the first steps toward a well-defined measurement theory in quantum field theory. Indeed, given that QFT is likely the most successful physical theory ever developed, any satisfactory solution for the Measurement Problem must also be an acceptable solution for measurement in QFT. Therefore, it is worrisome to note that researchers on the Measurement Problem do not usually deal with QFT (and vice versa).
Most scientists who use quantum mechanics in their daily work focus on its practical applications rather than its interpretation. Indeed, it makes no difference whether we interpret the result of an experiment as the quantum state collapsing into a single branch or experiencing a branch that cuts off communication with all others. Such a subtle distinction belongs to defining a preferred worldview. It is even less consequential than choosing between Newtonian and Hamiltonian formulations of classical mechanics, which at least differ in their practical use. The exception concerns objective collapse theories, which propose new physics beyond conventional quantum mechanics. In any case, it makes sense to keep challenging all interpretations and theories to prune the weak and reach complete consistency on the remaining, even if we never get a single consensual description of quantum mechanics.